OpenStudy (anonymous):

how do u find the focus, directrix and focus diameter of a parabola with the equation of x^2=y

OpenStudy (anonymous):

the general form of equation is x^2=4ay, then the focus is (0,a) , directrix is y=-a

OpenStudy (amistre64):

the general from of a parabola is x^2 = (4a)y ; where a is the distance from the vertex to the foci in one direction; and the distance between the vertex and the directix in the other direction

OpenStudy (amistre64):

x^2 = y has no extra numbers to shift it around; so we say it is centered at the origin; with a vertex of (0,0)

OpenStudy (amistre64):

x^2 is always a positive number; so the parabola opens upwards

OpenStudy (anonymous):

k, but if i am trying to graph it the focus would be (0,1) ?

OpenStudy (amistre64):

4ay = y ; /y 4a = 1 ; /4 a = 1/4

OpenStudy (anonymous):

x^2=4(1/4)*y so the coordinates of the focus is (0,1/4) and y=-1/4 is the equation of directrix...

OpenStudy (anonymous):

its ust confusing because i dont have the a and y in the original equation

OpenStudy (amistre64):

it best to commit the general equation to memory if you are wanting to pursue a career in the maths. x^2 = 4ay; is the general x^2 = y; is the given which means that 4ay = y and then solve for 'a'

OpenStudy (anonymous):

k im following ya

OpenStudy (amistre64):

that is the process I posted above; since 'a' is the determing part of the question

OpenStudy (amistre64):

4ay = y ; divide out the y 4a = 1 ; divide out the 4 a = 1/4 ; and this is the distance from the vertex to the foci; and the distance from vertex to the directix

OpenStudy (amistre64):

the foci is in the direction of the upward curve; and the directix sits behind the vertex

OpenStudy (amistre64):

vertex = (0,0) ; and graph opens up; so the y component increases by |a| foci = (0,1/4) directix is the other side of it; y = 0-1/4; y=-1/4